OpenStudy (anonymous):

Change the given equation into slope-intercept form. 5x – 2y = 10

OpenStudy (anonymous):

they are asking you to solve for y

OpenStudy (anonymous):

\[5x-2y=10\] \[5x-10=2y\] \[\frac{5}{2}x-5=y\]